ABC Corp. allocates overhead to production on the basis of direct labor costs. If Deltan's total estimated overhead is $450,000 and estimated direct labor cost is $180,000, determine the amount of overhead to be allocated to finished goods inventory. There is $20,000 of total direct labor cost in the jobs in the finished goods inventory. $20,000 $90,000 $70,000 $50,000
Answer #1

Overhead allocation rate = Estimated overheads/Direct labor Cost

= 450,000/180,000

= 2.5 per Dollar of labor cost

Hence, amount of overhead applied to Finished Goods Inventory = 20,000*2.5 = $50,000

Hence. the answer is $50,000